TheTiredCat:

[Half-Life Unit] 32P has a half-life of 14.3 days. How much of a 100.0 gram sample will remain after three half-lives?

Vocaloid:

Late response but here’s the formula for half life decay N(t) = N0 * (1/2) ^ (t/h) Where N(t) is the final amount, N0 is the initial amount, t is the total time and h is the number of half lives. Since you’re told 3 half lives have passed, t/h = 3. N0 is the initial mass 100g. Plug in and evaluate N(t).

TheTiredCat:

@vocaloid wrote:
Late response but here’s the formula for half life decay N(t) = N0 * (1/2) ^ (t/h) Where N(t) is the final amount, N0 is the initial amount, t is the total time and h is the number of half lives. Since you’re told 3 half lives have passed, t/h = 3. N0 is the initial mass 100g. Plug in and evaluate N(t).
tyty TvT I've needed this--